How do I solve this by using substitution? y=2x+3 and y=7x-2
Answer by EdenWolf(517)

We know what y equals, so plug 2x+3 into y in the second equation and solve for x first.

2x+3=7x-2
2x+5=7x
5=5x
x=1

So we know what x is, so now we find y. Plug the value of x, or 1, into any of the two equations. (Let's use the first one.)

y=2*1+3
y=2+3
y=5

So we know that x = 1, and y = 5.